# Computational techniques for sheaf cohomology of locally profinite sets

### v1

@@ 5. Cocycles as pointwise limits of coboundaries
$x^{\infty}_{i}=\begin{cases}x_{i},&i\not\in B,\\ \infty,&i\in B.\end{cases}$
@@
These choices assemble into a function
@@
is continuous. This in turn, holds if and only if the support of $\varphi_{\vec{f},B}^{\vec{y}}$ is a clopen subset of $\prod_{i\not\in B}\alpha(\kappa_{i})$. But continuity of $f_{B}$ guarantees that
@@
(Here, by “even” we mean “of the form $\omega\cdot\lambda+2k$ for $\lambda$ limit, $k<\omega$”. This can be replaced with any sequence of unbounded, co-unbounded subsets of the $\kappa_{i}$.) Notice this is well-defined even if any subset of the $m_{i}$ are set to $\infty$; thus $\varphi$ is continuous. Moreover, if not every $m_{i}$ is $\infty$, the resulting map is mod-finite constant in the $\beta_{0}$-direction, and is thus a coboundary. So, modulo Theorem C, we need only show that the map
@@ 6. Further work and conclusion
  • $H^{2}(X(\aleph_{0},\aleph_{0},\aleph_{0},\aleph_{1})^{-},\underline{\mathbb{Z}_{2}}=0$? $H^{2}(X(\aleph_{0},\aleph_{0},\aleph_{0},\aleph_{1})^{-},\underline{\mathbb{Z}_{2}}\neq 0$?

### v2

@@ 5. Cocycles as pointwise limits of coboundaries
$x^{\infty}_{i}=\begin{cases}x_{i},&i\in B,\\ \infty,&i\not\in B.\end{cases}$
@@
The $\varphi_{\vec{f},B,\vec{x}}^{\vec{y}}$ then assemble into a function
@@
is continuous. This in turn holds if and only if the support of $\varphi_{\vec{f},B}^{\vec{y}}$ is a clopen subset of $\prod_{i\not\in B}\alpha(\kappa_{i})$. But continuity of $f_{B}$ guarantees that
@@
(Here, by “even” we mean “of the form $\lambda+2k$ for $\lambda$ a limit ordinal, $k<\omega$”. This can be replaced with any sequence of unbounded, co-unbounded subsets of the $\kappa_{i}$.) Notice this is well-defined even if any subset of the $m_{i}$ are set to $\infty$; thus $\varphi$ is continuous. Moreover, if not every $m_{i}$ is $\infty$, the resulting map is mod-finite constant in the $\beta_{0}$-direction, and is thus a coboundary. So, modulo Theorem C, we need only show that the map
@@ 6. Further work and conclusion
  • $H^{2}(X(\aleph_{0},\aleph_{0},\aleph_{0},\aleph_{1})^{-},\underline{\mathbb{Z}_{2}})=0$? $H^{2}(X(\aleph_{0},\aleph_{0},\aleph_{0},\aleph_{1})^{-},\underline{\mathbb{Z}_{2}})\neq 0$?